If the angle between vector sand b is an acute angle  then the difference a_b is

Grade:12

1 Answers

Arun
5 years ago
Dear student
 
| a-b |= sqrt (a² +b² - 2 ab cos @)
As @ is acute angle
Henceif @>0 but @
| a-b | > |a| - |b|
& if @ = 0°
|a-b| = |a| - |b|